OpenStudy (anonymous):

a triangle has side lengths of 5,8, and 12. determine whether the triangle is right, acute or obtuse.

OpenStudy (smarty):

Acute

OpenStudy (accessdenied):

This triangle is obtuse. Consider the triangle with side lengths \(a \text{, } b \text{, and } c\) and opposite angles \(A \text{, } B \text{, and } C\). The Law of Cosines applies: \(c^2 = a^2 + b^2 - 2ab \cos C\) Two cases (the third is already covered by Pythagorean Theorem): \(\quad c^2 < a^2 + b^2\) if \(\neg 2ab \cos C < 0\) \(\quad c^2 > a^2 + b^2\) if \(\neg 2ab \cos C > 0\) Let's try to express this in terms of the angle only. We can divide off the \(\neg 2ab\) to simplify into terms of \(\cos C\) (flipping the sign, of course). \(\quad c^2 < a^2 + b^2\) if \(\cos C > 0\) \(\quad c^2 > a^2 + b^2\) if \(\cos C < 0\) C is the angle opposite of the c-value. Consider the possible values of C, in terms of our triangle: \(0 < C < 180\). We can then determine the values of C for which \(\cos C\) fits each of the two conditions: \(\quad \cos C > 0\) when \(0 < C < 90\) \(\quad \cos C < 0\) when \(90 < C < 180\) Using this information, we can conclude: . When C is an acute angle, then \(\cos C\) is greater than zero; therefore, \(c^2 < a^2 + b^2\) in this case. . When C is an obtuse angle, then \(\cos C\) is less than zero; therefore, \(c^2 > a^2 + b^2\) in this case. So, now let \(a = 5\), \(b = 8\), and \(c = 12\) \( \quad 12^2 \ \_ \ 5^2 + 8^2 \\ \quad 144 \ \_ \ 25 + 64 \\ \quad 144 \ \_ \ 89 \\ \quad 144 > 89 \) Therefore, this is \(c^2 > a^2 + b^2\); therefore, the triangle is obtuse.